Rewrite the equation 2x-3y=-6 as a function of x
Please help me on this to solve it.
A function notation problem
Thanks

Answer by stanbon(75887)

Rewrite the equation 2x-3y=-6 as a function of x
-------------
3y=2x+6
y=(2/3)x+2
f(x)=(2/3)x+2
--------
Cheers,
Stan H.